OpenStudy (anonymous):

Suppose the area of a rectangle is 7.9 in2 and the length is 2.1 in. longer than the width. Find the length and width in inches of the rectangle.

OpenStudy (anonymous):

So area=7.9 You know that area is width*height. You are also given L=W+2.1 So 7.9=W*(W+2.1) Distribute that to get 7.9=W^2+2.1W From there you move 7.9 to the other side to get a quadratic, so W^2+2.1W-7.9=0 Use a graphing calculator to solve from there. W=-4.05042, 1.95042 We can eliminate the negative answer as width cannot be negative. So Width=1.95042 Go back to L=W+2.1. Substitute W in, so L=1.95042+2.1 L=4.05042 W=1.95042

OpenStudy (anonymous):

This one does not give exact values Best I could find was that the width = 1.95 and length = 4.05

OpenStudy (anonymous):

x(x-2.1) = 7.9 x^2 -2.1 x - 7.9 = 0 solve this using formulas gives 4.05 or -1.95 x must be positive so length = 4.o5 and width = 4.05 - 2.1 = 1.95